2. A particle with a charge of 5nC has a distance of 0.5m away from a charge of 9.5nC. What is

its electric potential energy?

3. Find the electric potential at a distance of 8cm from a 7.58pC charge.

4. What is the electric potential given 0.58m away from a 18uC charge?

5. Calculate the amount of electric potential energy possessed by 2.52x10^-10C at a point

where the electric potential is 4.7 Volts.

6. The electric potential at a distance of 0.92m from a point charge is -8.41x10^5 V. what is the

quantity of the point charge?


1
Expert's answer
2022-03-06T18:06:30-0500

2.

"U=\frac{kq_1q_2}r=855~nJ,"

3.

"\varphi=\frac{kq}r=0.85~V,"

4.

"\varphi=\frac{kq}r=279~kV,"

5.

"U=q\varphi=1.2~nJ,"

6.

"q=\frac{\varphi r}{k}=-86~\mu C."